OpenStudy (help!!!!):

find 3 consecutive odd integers such that the square of the first decreased by the square of the third is equal to 10 less than twice the square of the second

OpenStudy (help!!!!):

\[x^2-(x+3)^2=2(x+2)^2-10\]

OpenStudy (anonymous):

distribute the brackets first

OpenStudy (help!!!!):

This is what I got so far, but I know something is wrong

OpenStudy (anonymous):

the intengers should look like these x x+2 x+4..........because you need odd intengers

OpenStudy (help!!!!):

FYI I translated the problem as an equation but something is wrong. HELP!!!!

OpenStudy (mertsj):

x x+2 x+4 \[x^2-(x+4)^2=(x+2)^2-10\]

OpenStudy (anonymous):

instead of x+3 you should have x+4, just like @Mertsj wrote

OpenStudy (help!!!!):

lol I got it

OpenStudy (help!!!!):

Thanks for the help guys!!

OpenStudy (anonymous):

ok. np :)

OpenStudy (mertsj):

yw